OpenStudy (anonymous):

Find the quotient 27m^5 ------ 3m^3

OpenStudy (owlcoffee):

You can "split" them into the producto of two more simple fractions in order to operate them more easily: \[\frac{ 27m^5 }{ 3m^3 }\] the fraction can be written as: \[(\frac{ 27 }{ 3 })(\frac{ m^5 }{ m^3 })\]

OpenStudy (anonymous):

can u help me with solving the problem

OpenStudy (phebe):

\(WELCOME~ TO~ OPENSTUDY!\)

OpenStudy (owlcoffee):

Yes, first, let's center on the left fraction. what is 27 divided by 3?

OpenStudy (anonymous):

9

OpenStudy (owlcoffee):

good, and we transcribe that: \[(9)(\frac{ m^5 }{ m^3 })\] Now, we have to use an exponential property in order to solve the other, do you know which one it is?

OpenStudy (anonymous):

no